# PET-SQL: A Prompt-enhanced Two-stage Text-to-SQL Framework with Cross-consistency

### v3

@@ 4.3 The Effect of Prompts
Datasets Prompting Codellama-34B InternLM-70B SQLCoder-34B
Dev $CR_{p}$ 0.685 0.701 0.562
$RE_{p}$ 0.738 0.707 0.657
Test $CR_{p}$ 0.694 0.649 0.631
$OD_{p}$ 0.609 0.673 0.579
$RE_{p}$ 0.717 0.699 0.678
Ablation of $RE_{p}$
Since there are three modifications in $RE_{p}$ compared to $OD_{p}$, that is, optimization rule (OR), cell values references (CV), and foreign key declarations (FK), we perform the ablation experiments on the Spider dev set to provide insights into these components of the proposed objective function specifically. Table 4 shows the ablation results of the $RE_{p}$ on three non-OpenAI LLMs under the zero-shot setting. It is observed that the exclusion of the CV consistently exerts the most significant impact on performance across InternLM and SQLCoder, resulting in about 2.4% and 6.4% decreases in EX. The absence of the OR also demonstrates a notable impact on model performance (about -3%). Interestingly, the FK component exhibits a comparatively smaller influence on $RE_{p}$, even achieving slight improvement compared to the full version. These findings suggest that the CV has the greatest significance, followed by the OR and FK components. This indicates that ensuring the presence of sampled cell values is crucial for LLM generation, and with these references, LLM can better navigate the database and formulate accurate queries. This reduces ambiguity and improves the efficiency of the generated SQL queries.
Table 4: The ablation of the $RE_{P}$ on Spider dev set (0-shot)
CodeLlama-34B InternLM-70B SQLCoder-34B
$RE_{p}$ 73.80% 70.70% 65.70%
        w/o OR 71.50% 69.90% 64.60%
        w/o CV 71.30% 69.10% 61.70%
        w/o FK 66.30% 71.60% 62.30%
4.4 The Effect of Schema Linking
The linked schema is parsed from the PreSQL, which is generated by GPT4-0613 with full schema information. To validate its effect, we introduce the table recall metrics and compare the EX of FinSQLs generated by the LLMs based on simplified and unsimplified schemas, separately.
The table recall metrics are two-fold, and are defined as
  • Totally Equal: $R_{e}={\sum_{i=1}^{N}\mathbf{1}_{e}}/{N}$, where $N$ is the number of instances in the test set. $\mathbf{1}_{e}$ is an indicator function, which returns 1 if the linked tables are exactly the tables that appeared in ground-truth SQL (GT tables) and 0 otherwise. $R_{e}=1$ is the ideal schema linking result, which means all the question-related tables are recalled, and there is no verbose information.
Table 4.4 shows the schema linking (SL for short) recall of GPT4 (i.e., PreSQL). It can be seen that the recall of table linking reaches 94% while the EX of PreSQL is only 85.2%, and the limitations of LLMs may cause the gap. Prompt simplification with linked tables enables LLM to potentially correct the wrong cases in PreSQL. In Table 4.4, we further summarize the average tables mentioned in the prompt and the ground truth (i.e., the average of GT tables). In the full schema (without SL), the average number of tables is 4.89, while in the simplified schema (with SL), it is 1.60 (close to GT). This means that SL significantly simplifies useless table information in prompts.
Table 5: The metrics of SL with GPT4
@@ 4.4 The Effect of Schema Linking
GPT4 0.94 0.98
@@
With the linked tables, we further simplify the prompt and then evaluate the performance of LLMs on the test set with the 9-shot setting. The results are demonstrated in Table 7. The simplified prompt has improved performance across all LLMs, especially for CodeLlama, which validates the effectiveness of the PreSQL-based schema linking.
Table 7: The comparison of the LLMs performance with and without SL (9-shot)
Codellama-34B InternLM-70B SQLCoder-34B GPT4
w/o SL 74.4% 71.4% 74.1% 85.2%
w/ SL 78.2% 75.4% 76.9% 85.5%
4.5 The Effect of Cross-consistency
We evaluate the cross-consistency (CC for short) of the five LLMs mentioned above. In Table 8, we report its performance on dev and test sets. Considering the token cost, we give up the SQL generation by GPT4 on the dev set, and the best result from the remaining LLMs is chosen as the baseline (CodeLlama-34B). For comparison, we also perform self-consistency with CodeLlama-34B (i.e., calling it 5 times) under $temperature=0.5$. On the dev set, compared to the baseline, the gain of self-consistency is trivial, which is only 0.8%, while the cross-consistency with naive voting achieves 7.9% improvements. This proves that cross-consistency is much better than self-consistency. On the test set, it can seen that the fine-grained voting strategy achieves remarkable improvements over the baseline, with 2.1% boosting. Even for the naive version, the improvements are also promising (+1.2%). Compared to naive voting, fine-grained voting can further exploit the potential of LLMs and significantly mitigate the voting bias. For information, self-consistency of five SQLs generated by GPT4 only achieves 0.4% improvements in DAIL-SQL [4]. These findings demonstrate that, despite its power, GPT-4 may generate errors that other LLMs would not. Therefore, introducing several LLMs to complement each other (i.e., cross-consistency) is necessary, which can diversify the results and greatly improve the overall EX.
Table 8: The performance of cross consistency (9-shot)
Methods Dev(w/o GPT4) Test(w/ GPT4)
Baseline (best single LLM) 75.3% 85.5%
Self-consistency 76.1% -
CC(Naive voting) 82.2% 86.7%
CC(Fine-grained voting) - 87.6%
5 Conclusion
This paper presents the PET-SQL framework, which aims to address challenges in Text2SQL tasks by enhancing the prompt and leveraging cross-consistency across LLMs. Our approach achieves 87.6% execution accuracy on the Spider leaderboard. We also propose a PreSQL-based schema linking method to simplify prompt information, improving the efficiency and accuracy of LLMs in generating SQL queries. Overall, the PET-SQL framework demonstrates promising results and opens avenues for further advancements in Text2SQL tasks.

### v4

@@ A.6 Further analysis of the proposed prompt $RE_{p}$
Datasets Prompting Codellama-34B SQLCoder-34B InternLM-70B
Bird-dev $CR_{p}$ 0.333 0.266 0.265
$RE_{p}$ 0.355 0.345 0.308
Spider-dev $CR_{p}$ 0.685 0.562 0.701
$RE_{p}$ 0.738 0.657 0.707
Spider-test $CR_{p}$ 0.694 0.631 0.649
$OD_{p}$ 0.609 0.579 0.673
$RE_{p}$ 0.717 0.678 0.699
A.7 Further analysis of the PreSQL-based schema linking
Table 12: The metrics of SL with GPT-4
@@ A.7 Further analysis of the PreSQL-based schema linking
GPT-4 0.94 0.98
Table 12 shows the schema linking (SL for short) recall of GPT-4 (i.e., PreSQL). It can be seen that the recall of table linking reaches 94% while the EX of PreSQL is only 85.2%, and the limitations of LLMs may cause the gap. Prompt simplification with linked tables enables LLM to correct the wrong PreSQL cases.
@@
In Table 13, we further summarize the average tables mentioned in the prompt and the ground truth (i.e., the average of GT tables). In the full schema (without SL), the average number of tables is 4.89, while in the simplified schema (with SL), it is 1.60 (close to GT). This means that SL significantly simplifies useless table information in prompts.
A.8 Further analysis of the cross consistency
GPT-4’s pivotal role is uncovered in enhancing the final outcomes. We conduct experiments to gauge its significance. Table 14 demonstrates the importance of GPT-4 within the Spider-test set: Specifically, utilizing GPT-4 alone yielded an accuracy of 85.50%, showcasing its substantial impact. Conversely, excluding GPT-4 from the other four models used for CC led to a noticeable decline in performance, resulting in an accuracy of 79.80%. Replacing GPT-4 with a weaker model (InternLM2-20B) within the CC framework further exacerbated this decrement, yielding an accuracy of 79.08%. However, employing all five models within the CC framework resulted in a notable enhancement, achieving an accuracy of 87.60%. These findings underscore GPT-4’s indispensable contribution to the overall performance, illustrating its pivotal role in enhancing CC.
Table 14: The importance of GPT-4 on Spider-test set
Models EX
GPT-4 Only 85.50%
The other four models w/o GPT-4 for CC 79.80%
Rep. GPT-4 with a weak model for CC 79.08%
All the five models for CC 87.60%